The average mass of 50 pumpkins is 2.1kg . If one pumpkin is added the average mass is 2.2kg . What is the mass of the extra pumpkin
Found 2 solutions by Theo, greenestamps:
Answer by Theo(13342)

total mass is 50 * 2.1 = 105
let x equal the weight of the added pumpkin.
you get (105 + x) / 51 = 2.2
multiply both sides of this equation by 51 to get:
105 + x = 51 * 2.2 = 112.2
subtract 105 from both sides of this equation to get:
x = 112.2 - 105 = 7.2
the mass of the extra pumpkin is 7.2 kilograms.
that's your solution.
50 * 2.1 = 105 + 7.2 = 112.2 / 51` = 2.2

Answer by greenestamps(13208)



You can solve this kind of problem about averages by thinking about how the weight of each pumpkin compares to the new average.

The weight of the added pumpkin has to be above the new average by the amount that the total weight of the original 50 pumpkins is short of the new average.

Each of the original 50 pumpkins is short of the new average by (2.2-2.1) = 0.1kg. Overall, the total weight of the original 50 pumpkins is 50(0.1) = 5kg under what is required to make the new average.

So the weight of the added pumpkin has to be 5kg above the new average.

ANSWER: 2.2kg + 5kg = 7.2kg